OpenStudy (kkutie7):

Not really sure how to write the first four terms. I'm guessing f, f',f",f"'. Use known series to expand sqrt(a+R) in terms of R/a when |R| is much smaller than a. Display the first 4 non-zero terms.

OpenStudy (kkutie7):

\[\sqrt{a+R}\rightarrow \sqrt{a}\sqrt{1+\frac{R}{a}}\]

OpenStudy (anonymous):

derivative i guess what is the variable? \(a\)? what are you expanding about?

OpenStudy (anonymous):

you got some goofy questions tonight for sure

OpenStudy (kkutie7):

Honestly no idea. Yeah my professor thinks this is good practice for our exam tomorrow.

OpenStudy (anonymous):

actually now that i read the question, probably not successive derivatives it says "use known expansion"

OpenStudy (kkutie7):

oh so taylor series then...

OpenStudy (kkutie7):

so like the series for (1+x)^k

OpenStudy (anonymous):

maybe \[\sqrt{1+x}\] if \(R\) is much smaller than \(x\) this is expanding about 0 maybe someone else has a better idea

OpenStudy (anonymous):

yeah like that, for \(k=\frac{1}{2}\)

OpenStudy (kkutie7):

but what about sqrt(a) do I just kind of bring it along for the ride or?

OpenStudy (anonymous):

maybe take the expansion for \[\sqrt{1+x}\] replace \(x\) by \(\frac{R}{a}\) then multiply by \(\sqrt{a}\)

OpenStudy (anonymous):

like i said (maybe)

OpenStudy (kkutie7):

alright I'll see if that works out

OpenStudy (anonymous):

lol good luck!